OpenStudy (anonymous):

A certain battery is rated at 80 ampere hour. How many coulomb of charge can this battery supply

OpenStudy (ash2326):

It can supply=80*3600 Coulombs of charge

OpenStudy (jamesj):

because Charge = (Current) x (time in seconds) and 1 hour = 60 minutes = 60^2 seconds = 3,600 seconds

OpenStudy (shayaan_mustafa):

Ampere hour is the unit of electrical charge. Use formula of current, which is, I=q/t and find out "q" convert hours into sec. t=1hrs=1*3600=3600s I=80A q=It q=80*3600 q=288000Coulombs/seconds Regards, Read my Profile. :D